OpenStudy (anonymous):

a = a is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (shane_b):

Reflexive property of equality

OpenStudy (shane_b):

Here, if you have similar questions: http://library.thinkquest.org/2647/algebra/post.htm

OpenStudy (lgbasallote):

symmetric property is a+ b = a + b transitive property is if a = b and b = c then a = c reflexive property is a = a commutative property of multiplication is ab = ba